Find the total surface area in square meters of a rectuangular solid whose length is 7 meters, width is 6 meters, and depth is 3 meters.
A. 32 B. 162 C. 81 D. 126 E. None of these.

Answer by wasim(2)

total surface area of cuboid= 2(length*width+width*depth+length*depth)
2(7*6+6*3+7*3)=2(42+18+21)=162 square meters
Answer (B)